(1 point) Find the most general real-valued solution to the linear system of differential equations x1 () = C1 + C2 x2(t)I AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

Given el - 2 20 - 1 -4 ut O E 4 Find A values Now of ei gen >> (A-I ) - 0 ㅂ - - 0 -A- (4+4) (0+2) + 1 = 0 - 2²+67+ g = 0 T (+Sove (A +31) * = 0 Augmented matre lo R2 Re&R I 15 0 0 21 (t) = X(t) x, (t) xalt) - xlt) 1 ut M2 It) = 1 A eigen ve cor 11Now :):):) T e,te = (e, te) = 1 al का eite, - - 1 e = rei ei е : l. - loe, put e, e (:) Xt) с е *** (1) * 03 fee** () +33(0))